George Washington was born 11 years before Thomas Jefferson. In 1770 Washington's age was 3 more than 7 times the age of Jefferson in 1748. What was the sum of the two men's ages in 1750?
Answer by god2012(113)

Solution :
Let G be the age of George Washington and T be the age of Thomas in the year 1748.
George was born 11 years before Thomas.
Therefore , G = T + 11 always.
In 1770, let the age of George be X and that of Thomas be Y,
Now, X = G+22 , Y = T + 22
But X-Y = 11 always.
X = 3 + 7T = G + 22 = T+11+22 = T+33
6T = 30 OR T = 5 years. and G = 5 + 11 16 years in 1748.
In 1750, T = 7, G = 18
Therefore , Sum of G and T in 1750 is 25 years